f(x+y)=f(x).f(y)andf(5)=2 f'(0)=3 then f'(5)=? for all value of'x'and'y'

Grade:

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Praveen

Put x=y=0 the f(0)=f2(0)

f(0)=0 or 1

f'(0)=lim h->0 (f(0+h)-f(0))/h= (f(0)f(h)-f(0))/h= f(0)*(f(h)-1)/h

now f(0) cant be zero as f'(0) is 3

f(0)=1

lim h->0 (f(h)-1)/h = f'(0)=3  .............(1)

f'(5)=lim h->0 (f(5+h)-f(5))/h= (f(5)f(h)-f(5))/h= f(5)*(f(h)-1)/h

f'(5)=f(5)*3=2*3=6     using (1)

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

Dimple Ravi Naik
3 years ago
 
f(x+y)=f(x)f(y)
f(0+0)=f(0).f(0)
Therefore, f(0)=1,0
f`(x+y)=f'(x)f(y)+f'(y)f(x)
f(0+5)=f'(0)f(5)+f'(5)f(0)
 
 
Using f(0)=0
f'(5)=6
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Put x=y=0 the f(0)=f2(0)
f(0)=0 or 1
f'(0)=lim h->0 (f(0+h)-f(0))/h= (f(0)f(h)-f(0))/h= f(0)*(f(h)-1)/h
now f(0) cant be zero as f'(0) is 3
f(0)=1
lim h->0 (f(h)-1)/h = f'(0)=3 .............(1)
f'(5)=lim h->0 (f(5+h)-f(5))/h= (f(5)f(h)-f(5))/h= f(5)*(f(h)-1)/h
f'(5)=f(5)*3=2*3=6 using (1)

Thanks and Regards